OpenStudy (anonymous):

How many ways you can select two numbers till 50,which will their product as a multiple of 3?

OpenStudy (keenan):

is that between 1and 50

OpenStudy (anonymous):

yue.

OpenStudy (anonymous):

it has to be keenan mutibles of 50 not divisors

OpenStudy (keenan):

ok

OpenStudy (anonymous):

for example - 1*3 =3 5*3=15 ...all in multiple of 3.

OpenStudy (anonymous):

the number should be betwwen 1 and 50

OpenStudy (anonymous):

15 times

OpenStudy (anonymous):

To get a two factor product out of 50 numbers ,we can choose the two number in 50C2ways. which we need to negate the pairs where in both numbers are not multiple of 3. Non multiple of 3 till 50 = 50 - 50/3 Approx equal to 34 Ans :- 50C2 - 34C2